# Fine-grained Hallucination Detection and Editing for Language Models

### 17 Jan 2024 v2

@@ Abstract
Large language models (LMs) are prone to generate diverse factually incorrect statements, which are widely called hallucinations. Current approaches predominantly focus on coarse-grained automatic hallucination detection or editing, overlooking nuanced error levels. In this paper, we propose a novel task—automatic fine-grained hallucination detection—and present a comprehensive taxonomy encompassing six hierarchically defined types of hallucination. To facilitate evaluation, we introduce a new benchmark that includes fine-grained human judgments on two LM outputs across various domains. Our analysis reveals that ChatGPT and Llama2-Chat exhibit hallucinations in 60% and 75% of their outputs, respectively, and a majority of these hallucinations fall into categories that have been underexplored. As an initial step to address this, we train Fava, a retrieval-augmented LM by carefully designing synthetic data generations to detect and correct fine-grained hallucinations. On our benchmark, our automatic and human evaluations show that Fava significantly outperforms ChatGPT on fine-grained hallucination detection by a large margin though a large room for future improvement still exists. Fava’s suggested edits also improve the factuality of LM-generated text, resulting in 5-10% FActScore improvements. [^1] Our code, data, and demo are available at https://fine-grained-hallucination.github.io/
@@ 1 Introduction
Large language models (LMs; Brown et al. 2020) can generate highly fluent and plausible text. However, these models are prone to produce hallucinations—factually incorrect generations—which often impedes their deployment in real-world applications Mallen et al. (2022). Prior work on hallucinations in natural language generation often assumes the presence of a specific reference text and focuses on studying faithfulness to the references Ji et al. (2023). On the contrary, escalating apprehensions have been articulated with respect to LM generations that are not grounded in any specific source text, but rather in world knowledge and facts Zhang et al. (2023); Huang et al. (2022).
@@
Figure 1: Overview of fine-grained detection system and the proposed taxonomy for hallucinations.
Such prevalence of hallucinations in LM outputs has motivated the development of automatic hallucination detection Min et al. (2023) or editing systems Gao et al. (2022); Chen et al. (2023). However, those existing systems tend to treat diverse types of factual errors uniformly, often simplifying them into binary categories such as “factual” or “not factual” with a focus on certain domains or specific types of errors. While entity-level contradictions (Figure 1 top) are usually evident and can be easily rectified with a single reference, LMs often fabricate entities that lack real-world existence (Figure 1 bottom). This variability underscores the need for a more fine-grained approach to hallucination detection for both model development and human verification, and we argue that a detection system should provide precise identification of hallucination sequences, discern different error types, and suggest potential refinements.
In this work, we introduce a new task, benchmark, and model to facilitate automatic fine-grained hallucination detection (Figure 1). We first introduce a new taxonomy of hallucinations and tasks (Section 3), and collect a high-quality benchmark to test models’ ability to provide fine-grained hallucination detection (Section 4). We then introduce a new model that is capable of providing span-level hallucination detection and editing (Section 5) as the first step towards building reliable fine-grained hallucination detection systems.
@@
Our taxonomy (Figure 2; Table 1) hierarchically classifies factual errors in LM generations into six categories. We specifically focus on hallucinations in information-seeking scenarios, where grounding to world knowledge matters. Based on this taxonomy, we propose a new task of fine-grained hallucination detection, in which a system needs to identify locations of factual errors in a response along with fine-grained hallucination types.
@@
We carefully annotate two widely-used LMs’ responses to diverse instruction-following and knowledge-intensive queries and construct the first fine-grained hallucination annotation benchmark. Our benchmark consists of about 400 responses of ChatGPT and Llama2-Chat 70B with span-level annotations of hallucination types based on our proposed taxonomy. Our benchmark reveals on average 1.9 and 3.4 hallucinations per response in ChatGPT and Llama2-Chat, respectively. It also shows that in addition to the widely-studied entity-level errors, other error types like unverifiable sentences or invented concepts make up over 60% of LM-generated hallucinations, highlighting the urgent need to study and detect such diverse types.
@@
Due to the expensive costs of annotations, creating training data for this new task is challenging and our preliminary experiments find proprietary LMs such as ChatGPT perform poorly in fine-grained detection. To overcome those challenges, we design an LM-based synthetic data generation process, resulting in 35k training instances. On this training data, we fine-tune a new retrieval-augmented LM, Fava, which can identify and mark hallucinations at span level using a unified syntax. We compare Fava with state-of-the-art LMs including ChatGPT on fine-grained hallucination detection and editing in LM outputs. Fava significantly outperforms ChatGPT with and without external knowledge by 22.1% on fine-grained hallucination detection and 14.2% on binary detection. Furthermore, Fava effectively fixes factual errors in diverse LM outputs—Fava editing improves FActScore Min et al. (2023) of Alapaca 7, 13B, and ChatGPT output by 4.4%, 9.3% and 3.3%, respectively.
@@ 2 Related Work
Hallucinations in natural language generation.
Taxonomy of hallucinations in certain downstream tasks such as summarization Pagnoni et al. (2021), text simplifications Devaraj et al. (2022), or knowledge-grounded dialogues Dziri et al. (2022) have been studied. Such prior taxonomies often assume the existence of specific source text and focus on faithfulness to the source Ji et al. (2023), which is qualitatively different from LMs’ factual hallucinations grounded in world knowledge. In addition, prior taxonomies include many task-specific categories that cannot be directly adapted to describe hallucinations in LM generations.
@@
Several recent works study methods to predict factuality by giving binary labels of a statement being factual or not Manakul et al. (2023); Min et al. (2023), or focusing on entity-level factual errors editing Gao et al. (2022); Chen et al. (2023). Recent surveys have attempted to categorize hallucinations in LM generations, but such categorizations often stretch the definition of hallucinations to broad error types in LM-generated text, which cannot be traced to specific spans in text (Huang et al., 2023; Zhang et al., 2023; Rawte et al., 2023). In this work, we propose a fine-grained taxonomy for fact-conflicting hallucinations in long-form text generation to improve factual consistency. While prior work often develops factuality verification systems on top of proprietary LMs Chern et al. (2023), we introduce a carefully designed training pipeline for a smaller yet more competitive fine-grained detection system. Unlike prior qualitative analysis on hallucinations in LM responses on certain knowledge-intensive queries (e.g., open-domain QA, biography generation; Liu et al., 2023b; Min et al., 2023), we annotate responses to diverse queries in multiple domains and reveal hallucination distributions vary across domains.
Figure 2: An overview of our fine-grained hallucination taxonomy.
Type Example ChatGPT Llama2
 Entity Lionel Andrés Messi was born on June 12 24, 1987. 46.9% 46.5%
 Relation Lionel Messi acquired was acquired by Paris Saint-Germain. 2.8% 2.5%
 Contradictory Messi has yet to gain captaincy for the Argentina national football team. 12.0% 14.1%
 Invented Messi is also famous for his discovery of the famous airplane kick technique. 4.6% 23.0%
 Subjective Lionel Messi is the best soccer player in the world. 15.0% 8.8%
 Unverifiable In his free time, Messi enjoys singing songs for his family. 18.7% 5.1%
Table 1: Examples of hallucinations with their percentage in our human-annotated data. We also show the percentage of those errors in ChatGPT and LLama2-Chat-70B outputs.
Fact verification.
While research on hallucination detection focuses on identifying errors in model-generated text, a related area of research focuses on identifying factual inaccuracies in human written claims (Bekoulis et al., 2021). Thorne et al. (2018) introduce a large scale dataset based on Wikipedia documents for the training and evaluation of fact verification systems. Other datasets in information-dense domains like fact-checking against scientific articles (Wadden et al., 2020) and news documents (Wang, 2017) have been proposed. Several systems have been developed for fact-checking in these settings (Thorne and Vlachos, 2018; Schuster et al., 2021; Nakov et al., 2021), but have not been tested for LM-generated text. While our work focuses on hallucination detection in LM generations, Fava can be adapted to fact-check human written claims as well.
@@ 3 Fine-grained Hallucination Detection
3.1 Focus and Definitions
In this work, we focus on open-ended text generation given information-seeking queries which often require factual knowledge. We define hallucinations as factually incorrect or unverifiable statements given external world knowledge. We operationalize world knowledge as documents from the web that are most relevant to the given statement/query according to a search algorithm. [^2] Errors in common sense, numerical, or logical reasoning are out of the scope of this work.
@@ 3.2 Hallucination Taxonomy
Based on our definition of hallucinations, we build a hierarchical taxonomy to categorize them into fine-grained error types. To develop a hallucination taxonomy for open-ended LM generations, we draw inspiration from error categories in prior task-specific taxonomies (Pagnoni et al., 2021; Devaraj et al., 2022) and introduce new categories to describe more complex errors surfacing in LM generations. We conducted a pilot annotation with 9 NLP experts to discuss and refine our taxonomy to ensure good coverage across diverse error types.
Figure 2 shows our taxonomies to classify LM hallucinations and Table 1 shows examples of each type of hallucination. We group hallucinations into two major categories: statements that contradict world knowledge (Type 1) and unverifiable statements (Types 2, 3, and 4).
@@
(1b)  Relation : Contradictory relation errors are another sub-category within contradictory statements where a semantic relationship described in a statement is incorrect. While (1a) often involves errors in nouns, (1b) involves verbs, prepositions, or adjectives, and can be fixed by correcting the incorrect relation.
(1c)  Contradictory : Contradictory statement errors refer to statements that entirely contradict relevant evidence from the web. These are cases where the full sentence is refuted by the information provided in a given reference.
(2)  Invented : Invented errors refer to statements with concepts that do not exist based on world knowledge. These are cases when the LM generates a non-existent or entirely fabricated entity that doesn’t occur in any relevant evidence. This does not include fictional characters in creative work such as books or movies.
(3)  Subjective : Subjective errors refer to an expression or proposition that lacks universal validity and is often influenced by personal beliefs, feelings, opinions, or biases and hence cannot be judged as factually correct. Specifically, this represents cases where the LM generates a statement without any factual proposition grounded in relevant evidence.
(4)  Unverifiable : Unverifiable errors refer to statements that contain factual propositions but cannot be grounded in world evidence. These are cases where the LM generates statements with facts, but none of the retrieved evidence from the web can directly support or contradict the fact (e.g., personal or private matters).
While  Entity  or  Relation  are often phrase-level and can be fixed by minimal editing erroneous phrases, other error types can be an entire sentence or part of a sentence and should be removed from a response to make it factual.
@@ 3.3 Tasks and Metrics
We introduce two new tasks of identifying and editing fine-grained factual errors in LM outputs. Given an input query $x$ and a corresponding LM output $y$, our tasks require systems to identify all of the factual errors. Each error $e$ consists of $(e^{text},e^{type})$, indicating the factually incorrect text spans and their error types among our taxonomies, respectively. We evaluate systems’ abilities concerning identifying fine-grained error types in the model-generated text $e^{type}$ (Task 1) and editing factual errors $e^{text}$ accordingly (Task 2).
Figure 3: Hallucination type distribution of human annotations across 3 datasets and two models.
@@
In the first fine-grained error detection task, the system is expected to identify fine-grained errors in an LM output. Due to the subjectivity of span-level annotations, for automatic evaluation, we evaluate systems’ abilities to detect whether an error type $t$ exists in a sentence $s_{i}\in y$. Given an output $y$ consisting of $L$ sentences, we assume the availability of gold error type annotations $e^{*t}_{i}\in\{\texttt{TRUE},\texttt{FALSE}\}$, which is a binary label of an error type $t$ existing in the $i$th sentence (TRUE) or not (FALSE). For each type, a system predicts $e^{t}_{i}$ and we evaluate precision and recall as follows:
$\text{Prec}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}==e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}==\texttt{TRUE}]}$ (1)
$\text{Recall}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}==e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{*t}_{i}==\texttt{TRUE}]}$ (2)
Precision indicates how many of the model’s predictions of an error type $t$ existing in the $i$th sentence is correct, while recall indicates how many of the error sentence TRUE is identified by the model. For final score, we compute the F1 scores averaged over six error types as follows:
@@
Fine-grained error detection can be simplified into a binary classification task, in which a system predicts if a sentence $s_{i}$ includes any factual errors or not. This is similar to FActScore Min et al. (2023), where we predict and evaluate factuality at the sentence level without extracting atomic facts.
@@
As shown in Table 1, some hallucinations can be fixed with minimal span-level edits, while others need to be removed or marked as unverifiable. In the editing task, the system is expected to suggest fine-grained edits to improve the factuality of the LM output. We evaluate a system’s ability to improve the factuality of given output $y$, by comparing scores estimated by off-the-shelf systems, such as FActScore or QAFactEval Fabbri et al. (2022). as follows:$f(\hat{y})-f(y)$, where $f$ indicates an evaluation function for factuality while $f(\hat{y})$ indicates edited output.
4 Benchmark
4.1 Human Annotation on Hallucinations
To facilitate model development in fine-grained hallucination detection and understand how prevalent those different types of hallucinations are, we carefully annotate the first fine-grained hallucination detection benchmark. Our benchmark consists of multi-way fine-grained annotations on two LM responses to queries in multiple domains.

### 21 Feb 2024 v3

@@ Abstract
Large language models (LMs) are prone to generate factual errors, which are often called hallucinations. In this paper, we introduce a comprehensive taxonomy of hallucinations and argue that hallucinations manifest in diverse forms, each requiring varying degrees of careful assessments to verify factuality. We propose a novel task of automatic fine-grained hallucination detection and construct a new evaluation benchmark, FavaBench, that includes about one thousand fine-grained human judgments on three LM outputs across various domains. Our analysis reveals that ChatGPT and Llama2-Chat (70B, 7B) exhibit diverse types of hallucinations in the majority of their outputs in information-seeking scenarios. We train Fava, a retrieval-augmented LM by carefully creating synthetic data to detect and correct fine-grained hallucinations. On our benchmark, our automatic and human evaluations show that Fava significantly outperforms ChatGPT and GPT-4 on fine-grained hallucination detection, and edits suggested by Fava improve the factuality of LM-generated text. [^1] Our code, data, and demo are available at https://fine-grained-hallucination.github.io/
Fine-grained Hallucination Detection and Editing for Language Models
[…]
@@ 1 Introduction
Large language models (LMs; Brown et al. 2020) can generate highly fluent and plausible text. However, these models are prone to produce factually incorrect or unverifiable statements, often called hallucinations. This impedes their deployment in real-world applications for information-seeking contexts Mallen et al. (2022). Prior work on hallucinations in natural language generation (NLG) often assumes the presence of a specific reference text and focuses on studying faithfulness to the references Ji et al. (2023). On the contrary, escalating apprehensions have been articulated concerning LM generations that are not grounded in any specific source text, but rather in world knowledge and facts Zhang et al. (2023).
@@
Figure 1: Overview of taxonomy for LM hallucinations and our model, Fava, a fine-grained detection system.
Figure 2: An overview of our fine-grained hallucination taxonomy. We identify 6 fine-grained types representing diverse hallucinations in LM-generated text.
Several recent work studies automatic hallucination detection Min et al. (2023) or editing outputs Gao et al. (2022) to address such LM hallucinations. These systems typically categorize hallucinations into simplistic binary distinctions like factual or not factual (Figure 1 top). We argue that hallucinations manifest in diverse forms, each requiring varying degrees of careful assessments to verify factuality. Entity-level contradictions are usually evident and can be easily rectified with a single reference. Conversely, errors involving fabricated entities (e.g., The Messi Diaries in Figure 1) demand thorough verification across multiple sources. This underscores the need for a more fine-grained approach to detect hallucination for model development and human verification.
@@
In this work, we propose automatic fine-grained hallucination detection (Figure 1 bottom), a new task requiring a system to provide precise identification of hallucination sequences, discern different types of hallucinations based on a taxonomy, and suggest refinements. We focus on hallucinations in information-seeking scenarios, where grounding to world knowledge matters. Our taxonomy (Figure 2) hierarchically classifies hallucinations in LM generations into six categories, based on careful pilot studies with NLP experts.
@@
We construct a new fine-grained hallucination benchmark, FavaBench, by carefully annotating approximately 1,000 responses of three widely used LMs (Llama2-Chat 7B, 70B and ChatGPT [^2] We use gpt-3.5-turbo-0301 throughout this work.) to diverse knowledge-intensive queries. Each response is annotated at the span level to identify hallucinations, including erroneous subspace, types, and potential refinements. Our analysis reveals all models include at least one hallucination in the majority of their responses (e.g., 70.2% in Llama2 7B and 59.8% in ChatGPT). In addition to the widely-studied entity-level errors, other error types like unverifiable sentences make up over 60% of LM-generated hallucinations, highlighting the urgent need to study and detect such diverse types.
@@
We introduce Fava, a new retrieval-augmented LM that can identify and mark hallucinations at the span level using a unified syntax (Figure 1, left bottom). Due to the annotation costs, we design an LM-based synthetic data generation process, and train Fava on the 35k resulting instances. We compare Fava with state-of-the-art LMs on fine-grained hallucination detection and editing in LM outputs. Fava significantly outperforms ChatGPT with and without external knowledge by 23.7% on fine-grained hallucination detection. Furthermore, Fava effectively corrects hallucinations in diverse LM outputs—Fava editing improves FActScore Min et al. (2023) of Alapaca 7, 13B, and ChatGPT by 4.4, 9.3 and 3.3%, respectively.
@@ 2 Related Work
Type Example ChatGPT Llama2-7b Llama2-70b
 Entity Lionel Andrés Messi was born on June 12 24, 1987. 42.7% 48.6% 55.7%
 Relation Lionel Messi acquired was acquired by Paris Saint-Germain. 4.7% 3.4% 2.8%
  Sentence Messi has yet to be captain for the Argentina football team. 18.9% 15.7% 19.9%
 Invented Messi is known for his famous airplane kick technique. 14.2% 17.4% 9.5%
 Subjective Lionel Messi is the best soccer player in the world. 9.9% 8.6% 5.8%
 Unverifiable In his free time, Messi enjoys singing songs for his family. 9.6% 6.3% 6.3%
Table 1: Distribution of different errors across ChatGPT, Llama2-Chat-7B and Llama2-Chat-70B outputs and examples of hallucinations for each type from our human-annotated data
Hallucinations in NLG.
Prior taxonomies proposed for summarization Pagnoni et al. (2021) or text simplifications Devaraj et al. (2022) often assume the existence of specific source text and focus on faithfulness to the source Ji et al. (2023). As this is qualitatively different from LMs’ factual hallucinations grounded in world knowledge, we extend prior work and build a new taxonomy for hallucinations in LM outputs.
@@
Several recent or concurrent studies propose methods to predict factuality by giving binary labels of a statement being factual or not Manakul et al. (2023); Min et al. (2023), or focusing on entity-level hallucination editing Gao et al. (2022); Chen et al. (2023). Recent surveys have attempted to categorize hallucinations in LM generations, often stretching the definition of hallucinations to broad error types in LM-generated text, which cannot be traced to specific spans in text (Huang et al., 2023; Zhang et al., 2023; Rawte et al., 2023). We propose a fine-grained taxonomy for hallucinations in long-form text generation for information-seeking context to improve factual consistency. While prior work often develops binary factuality verification systems on top of proprietary LMs Chern et al. (2023), we train a fine-grained detection system using our carefully designed data generation pipeline.
Fact verification for human-written claims.
While research on hallucination detection focuses on identifying errors in model-generated text, a related area of research focuses on identifying factual inaccuracies in human written claims (Bekoulis et al., 2021). Many large-scale datasets have been proposed for fact verification given Wikipedia Thorne et al. (2018), scientific articles (Wadden et al., 2020) or news documents (Wang, 2017). Several systems have been developed for fact-checking in these settings (Thorne and Vlachos, 2018; Schuster et al., 2021; Nakov et al., 2021), but have not been tested for LM-generated text. While our work focuses on hallucination detection in LM generations, Fava can be adapted to fact-check human written claims.
@@ 3 Fine-grained Hallucination Detection
Focus and definitions. We focus on open-ended text generation given information-seeking queries which often require factual knowledge. We define hallucinations as factual errors or unverified statements given external world knowledge. We operationalize world knowledge as documents from the web that are most relevant to the given statement/query according to a search algorithm. [^3] Errors in common sense, numerical, or logical reasoning are out of the scope of this work.
@@ 3.1 Hallucination Taxonomy
Based on our definition of hallucinations, we build a hierarchical taxonomy to categorize them into fine-grained error types. Inspired by prior task-specific taxonomies (Pagnoni et al., 2021; Devaraj et al., 2022), we introduce new categories to describe more complex errors surfacing in LM generations. We conducted a pilot annotation with nine NLP experts to ensure good coverage across diverse error types. Figure 2 shows our taxonomies to classify LM hallucinations and Table 1 shows examples of each type of hallucination. We group hallucinations into two major categories: statements that contradict world knowledge (Type 1) and unverifiable statements (Types 2, 3, and 4).
@@
(1b)  Relation : Contradictory relation errors are another sub-category within contradictory statements where a semantic relationship (e.g., verbs, prepositions, or adjectives) in a statement is incorrect.
(1c)   Sentence : Contradictory sentence errors refer to cases where a full statement entirely contradicts relevant evidence from the web.
(2)  Invented : Invented errors refer to statements where the LM generates an entirely fabricated entity that doesn’t exist based on world knowledge. Fictional entities in creative work aren’t included.
(3)  Subjective : Subjective errors refer to an expression that lacks universal validity and is often influenced by personal beliefs, opinions, or biases.
(4)  Unverifiable : This refers to statements where the LM output contains facts, but none of the retrieved evidence from the web can directly support or contradict the fact (e.g., personal matters). [^4] While  Invented and  Unverifiable are highly related,  Invented are the hallucinations where we can verify that some core entities or subjects of the sentences don’t exist.
While  Entity  or  Relation  are often phrase-level and can be fixed by minimal editing erroneous phrases, other error types can be an entire sentence or part of a sentence and should be removed from a response to make it precisely factual.
@@ 3.2 Tasks and Metrics
We introduce the two tasks of identifying and editing fine-grained factual errors in LM outputs. Given an input query $x$ and a corresponding LM output $y$, our tasks require systems to identify all of the factual errors in $y$. Each error $e$ consists of $(e^{text},e^{type})$, indicating the factually incorrect text spans and their error types among our taxonomies, respectively. We evaluate systems’ abilities concerning identifying fine-grained error types in the model-generated text $e^{type}$ (Task 1; Figure 1 < Entity >) and editing factual errors $e^{text}$ accordingly (Task 2; Figure 1 Barcelona $\rightarrow$ Miami).
Figure 3: Distribution of hallucination types in ChatGPT, Llama2-Chat-7B and Llama2-Chat-70B outputs across four datasets of diverse information-seeking queries based on our human annotated benchmark.
@@
In this task, the system is expected to identify fine-grained errors in an LM output. Due to the subjectivity of span-level annotations, for automatic evaluation, we evaluate systems’ abilities to detect whether an error type $t$ exists in a sentence $s_{i}\in y$. Given an output $y$ consisting of $L$ sentences, we assume the availability of ground-truth error type annotations $e^{*t}_{i}\in\{\texttt{TRUE},\texttt{FALSE}\}$, which is a binary label of an error type $t$ existing in the $i$th sentence (TRUE) or not (FALSE). Following the fact verification literature Thorne et al. (2018); Schuster et al. (2021); Feng et al. (2023) of computing precision and recall, we measure those metrics while we extend them for each error type. For each type, a system predicts $e^{t}_{i}$ and we evaluate precision and recall as follows:
$\text{Prec}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=\texttt{TRUE}]}$ (1)
$\text{Recall}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{*t}_{i}=\texttt{TRUE}]}$ (2)
Precision indicates the proportion of how many of the model’s predictions of an error type $t$ existing in the $i$th sentence is correct, while recall indicates how many of the error sentence TRUE is identified by the model. For final score, we compute the F1 scores averaged over six error types as follows:
@@
Fine-grained detection can be simplified into a binary classification task, in which a system predicts if a sentence $s_{i}$ includes any factual errors or not. [^5] This is similar to FActScore Min et al. (2023), while we predict and evaluate factuality at the sentence level without extracting atomic facts.
@@
Some hallucinations can be fixed with minimal span-level edits, while others need to be removed or marked as unverifiable. In the editing task, the system is expected to suggest fine-grained edits to improve the factuality of the LM output. We evaluate a system’s ability to improve the factuality of given output $y$, by comparing scores estimated by off-the-shelf systems, such as FActScore Min et al. (2023), as: $f(\hat{y})-f(y)$, where $f$ indicates an estimated factuality scores and $\hat{y}$ indicates edited output.
4 Benchmark: FavaBench
We create FavaBench to facilitate model development and evaluation in fine-grained hallucination detection and understand how prevalent those different types of hallucinations are. FavaBench consists of multi-way fine-grained annotations on three LM responses to queries in multiple domains.